OpenStudy (anonymous):

The resistors have the values Ra=2.00 ohm, Rb=9.00 ohm, Rc=1.00 ohm, Rd=1.00 ohm, Re=6.00 ohm, and the batteries, which have negiligible internal resistance, have emf's of Eu=3.50 V, Ev=9.50 V, Ew=1.50 V. Calculate the power dissipated in Re

OpenStudy (anonymous):

Calculate the current through Rd. Use a + sign for current flowing to the left in Rd. Calculate the power delivered by Ev.

OpenStudy (anonymous):

I know I need to add the left and right loops to get the currents, but I'm having trouble setting up the equations

OpenStudy (anonymous):

how do name the current in loop 1 and loop 2? I1 and I2 for instance?

OpenStudy (anonymous):

yeah, so I1 would be current through Re to the right, I2 would be current through Rc tupward and I3 would be current through Rd to the left

OpenStudy (anonymous):

No need to use three currents. Imagine you have I1, clockwise, passing through Ew, Ev, Rc,Re and Rb And I2 passing clockwise through Eu, Ra, Rd , Rc and Ev

OpenStudy (anonymous):

Ok not sure how to set that up then

OpenStudy (anonymous):

Is the equation for loop1 (forget about I2 now) I1(Rc+Re+Rb)+Ew+Ev=0?

OpenStudy (anonymous):

That looks correct to me. So would the second loop be I2(Ra+Rd+Rc)+Eu+Ev=0 or would it be different because its in a different direction?

OpenStudy (anonymous):

YEs except for Ev, whose sign has to be "minus" because I2 is hitting the battery by the negative part

OpenStudy (anonymous):

ok, so from those I can find the currents. To get the power dissipated there is it just current^2 * resistance? also not sure about power delivered

OpenStudy (anonymous):

No, no, bear with me, we have not finished yet the equations!

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

In the equation that I wrote, remember I told you to discard I2?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

What is the effect of I2 in Loop1? Isn't it -I2xRc

OpenStudy (anonymous):

yep but not sure what that gives you

OpenStudy (anonymous):

Just put -I2xRc in my equation and you have the right equation for loop1. Now you do the same for loop2. See the effect of I1 in loop2

OpenStudy (anonymous):

so you subtract that form loop one and add it to loop 2?

OpenStudy (anonymous):

The effect of I1 in Loop2 is -I1 x Rc. You get it? (I1 goes opposite to I2, therefore its effect on common parts to both loops will be negative)

OpenStudy (anonymous):

Put -I1 x Rc in your equation for loop2

OpenStudy (anonymous):

So it's opposite?

OpenStudy (anonymous):

Not sure where you're going with this

OpenStudy (anonymous):

Yep. Now you can tell me the complete equations for loop1 and 2

OpenStudy (anonymous):

I1(Rc+Re+Rb)+Ew+Ev-I2xRc=0 for loop 1?

OpenStudy (anonymous):

Yes, I give you the second. Check with mine: I1x(Rc+Re+Rb)-I2xRc=-Ew-Ev -I2xRc+I2x(Ra+Rd+Rc)=-Eu+Ev

OpenStudy (anonymous):

yep, looks correct. So from the currents then how do you get the power dissipated and delivered?

OpenStudy (anonymous):

The power is RI^2 and the current that passes through Re is I1, then you have it

OpenStudy (anonymous):

From the first equation I get I2 =2, does that seem correct?

OpenStudy (anonymous):

Woops, I mean second equation

OpenStudy (anonymous):

Since there is only one variable

OpenStudy (anonymous):

No, there are two variables. It is an equations system: Loop1: I1x(Rc+Re+Rb)-I2xRc=-Ew-Ev Loop2: -I2xRc+I2x(Ra+Rd+Rc)=-Eu+Ev

OpenStudy (anonymous):

Solve the system

OpenStudy (anonymous):

I don't see 2 variables in the second equation. Only I2

OpenStudy (anonymous):

Mistake of mine: The effect of I1 is -Rc x I1, then the second equation is: -I1xRc+I2x(Ra+Rd+Rc)=-Eu+Ev

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

Sorry for that

OpenStudy (anonymous):

no problem, give me a sec to work it out

OpenStudy (anonymous):

I end up with a negative number for I1, but that should be ok right?

OpenStudy (anonymous):

I get I1=-38/63

OpenStudy (anonymous):

yep same here

OpenStudy (anonymous):

And I2=85/63

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

You got it. What is the current through Re?

OpenStudy (anonymous):

-38/63

OpenStudy (anonymous):

That is, I1, then We=I1^2 x Re

OpenStudy (anonymous):

gives me 2888/1323 Jor 2.18 J

OpenStudy (anonymous):

We=(38/63)^2 x 6 in Watts

OpenStudy (anonymous):

oh woops yeah watts lol

OpenStudy (anonymous):

If I asked to calculate the power in Rc, what current would you use?

OpenStudy (anonymous):

I'm not sure. both?

OpenStudy (anonymous):

The current through Rd should just be I2, right?

OpenStudy (anonymous):

Yep, I1-I2 or I2-I1, is the same because for the power you use power two

OpenStudy (anonymous):

and through Rd is I2, true

OpenStudy (anonymous):

Great! so now for the power delivered by Ev?

OpenStudy (anonymous):

should be related to the current through Rc

OpenStudy (anonymous):

Maybe the way to solve is a bit fuzzy but see that it is like writing the equations for each loop but substracting the effect of the current of adjacent loops. like this you may solve any circuit no matter how complex.

OpenStudy (anonymous):

That makes sense. I think I need more practice

OpenStudy (anonymous):

Yes, but take into account that if I1-I2 is negative you can say that the battery is delivering power, otherwise we are charging it

OpenStudy (anonymous):

so Ev should just be (I1-I2)^2 * Rc

OpenStudy (anonymous):

No. It is (I2-I1)xEv (the battery delivers power for a current that is flowing out from the battery, that is whay I said I1-I2 has to be negative) Then I2-I1 has to be positive and the power is (I2-I1)xEv

OpenStudy (anonymous):

It gives me a number around 19 Watts

OpenStudy (anonymous):

Ah woops, was thinking about the wrong thing. Thank you, that worked and maeks more sense to me now. Time to go practice some more

OpenStudy (anonymous):

I leave you, I am going to hit the hay. Enjoy circuits and electronics. It is really amusing